# Probing false vacuum decay and bubble nucleation in a Rydberg atom array

### v1

@@ IV BCH expansion of AFM order dynamics
The BCH (nested–commutator) expansion provides a short-time series for Heisenberg-picture evolution. We use it to analyze the early-time dynamics of the AFM order parameter
@@ IV.A The vanishing odd orders
For any real initial state, all odd-order coefficients vanish in expectation. Let $\Theta$ denote time-reversal (complex conjugation in the $\hat{\sigma}^{z}$ basis). Since $\Theta\hat{H}\Theta^{-1}=\hat{H}$ and $\Theta\hat{M}\Theta^{-1}=\hat{M}$, any $|\psi\rangle$ that is real in this basis (e.g., Néel or PQG state), $\Theta|\psi\rangle=|\psi\rangle$, satisfies
@@ IV.B The second order
For the PQG state (defined as the ground state of $\hat{H}$ at $\Delta_{l}\to 0^{-}$), it is convenient to compute $\langle\hat{C}_{2}\rangle_{\mathrm{FV}}$ via a splitting of $\hat{H}$. Decompose $\hat{H}=\hat{H}_{0}+\hat{\delta}$ with
@@
Thus $\langle\hat{C}_{2}\rangle_{\mathrm{PQG}}$ is linear in $\Delta_{l}$; the expectation value $\big\langle\sum_{i}\hat{\sigma}_{i}^{x}\big\rangle_{\mathrm{PQG}}$ is taken with respect to the pre-quench ground (PQG) state, which does not depend on the coefficient $\Delta_{l}$ in the post-quench Hamiltonian.
The value $\langle\hat{C}_{2}\rangle_{\mathrm{PQG}}$ can be interpreted as follows: We treat the state $|\text{Néel}\rangle=|\downarrow\uparrow\dots\downarrow\uparrow\rangle$ as the background, on which an $L=1$ bubble at site $i$ can be created by the operator $\hat{\sigma}_{i}^{+}$ (for odd $i$) or $\hat{\sigma}_{i}^{-}$ (for even $i$). By mapping the spin operators to hard-core boson operators [10.1143/PTP.16.569, pitaevskii2016bose, fetter2003quantum], the expectation value $\langle\sum_{i}\hat{\sigma}_{i}^{x}\rangle_{\rm{PQG}}$ thus represents the condensate amplitude of the $L=1$ bubble.
The ratio of the second order
After approximating $V_{ij}$ as a nearest-neighbor interaction with strength $V$, and assuming $\Delta_{g}=V$ to map the system onto a transverse-field Ising model (TFIM), and then taking the thermodynamic limit of the ring with $N\to\infty$, we can obtain an analytical expression for the ratio of the two expectation values of $\hat{C}_{2}$, which reads
@@ IV.C The fourth order
When evaluating the expectation value of $\hat{C}_{4}$ under the PQG state, only terms with $\mathcal{D}$ acting outermost survive, $\langle\hat{C}_{4}^{(0)}\rangle_{\text{PQG}}=0$, such that
@@
For simplicity, we will take into account only the NN terms $V$ that dominate in long-range interaction $V_{ij}\propto|i-j|^{-6}$. In the region of small $\Delta_{l}$, the first-order $\mathcal{O}(\Delta_{l})$ contribution in Eq. (S16b) yields density-dressed single-bubble and pair-exchange blocks,
@@
When evaluating the expectation value of $\hat{C}_{4}$ under the Néel state, the lowest term in the $\Delta_{l}$-polynomial is the zeroth-order, $\mathcal{O}\left((\Delta_{l})^{0}\right)$. Using the definition of the Néel state, Eq. (S16a) yields:
@@ IV.D Higher orders and practical truncation
At higher orders, in the small-$\Delta_{l}$ limit one still finds $\langle\hat{C}_{2m}\rangle_{\mathrm{PQG}}=\mathcal{O}(\Delta_{l})$, indicating the slow decay of $|\text{PQG}\rangle$ in low $\Delta_{l}$ regime. At order $\hat{C}_{6}$ we encounter length-3 (tri-bubble) NN strings, e.g.

### v2

@@ IV BCH expansion of AFM order dynamics
The Baker-Campbell-Hausdorff (BCH) expansion provides a perturbative series for the short-time Heisenberg evolution. We employ this method to analyze the early-time dynamics of the AFM order parameter,
@@ IV.A The vanishing odd orders
For any initial state with real coefficients in the computational basis, all odd-order expansion coefficients vanish. Let $\Theta$ denote the time-reversal operator (complex conjugation in the $\hat{\sigma}^{z}$ basis). Since $\Theta\hat{H}\Theta^{-1}=\hat{H}$ and $\Theta\hat{M}\Theta^{-1}=\hat{M}$, any $|\psi\rangle$ that is real in this basis (e.g., Néel or PQG state), $\Theta|\psi\rangle=|\psi\rangle$, satisfies
@@ IV.B The second order
In contrast, for the PQG state (the ground state of $\hat{H}$ at $\Delta_{l}\to 0^{-}$), we decompose the Hamiltonian as $\hat{H}=\hat{H}_{0}+\hat{\delta}$:
@@
Thus $\langle\hat{C}_{2}\rangle_{\mathrm{PQG}}$ is explicitly proportional to the symmetry-breaking field $\Delta_{l}$.
The value $\langle\hat{C}_{2}\rangle_{\mathrm{PQG}}$ can be interpreted as follows: We treat the state $|\text{Néel}\rangle=|\downarrow\uparrow\dots\downarrow\uparrow\rangle$ as the background, on which an $L=1$ bubble at site $i$ can be created by the operator $\hat{\sigma}_{i}^{+}$ (for odd $i$) or $\hat{\sigma}_{i}^{-}$ (for even $i$). By mapping the spin operators to hard-core boson operators [52, 53], the expectation value $\langle\sum_{i}\hat{\sigma}_{i}^{x}\rangle_{\rm{PQG}}$ thus represents the condensate amplitude of the $L=1$ bubble.
After approximating $V_{i,j}$ as a nearest-neighbor interaction with strength $V$, and assuming $\Delta_{g}=V$ to map the system onto a transverse-field Ising model (TFIM), and then taking the thermodynamic limit with $N\to\infty$, we can obtain an analytical expression for the ratio of the two expectation values of $\hat{C}_{2}$, which reads
@@ IV.C The fourth order
When evaluating the expectation value of $\hat{C}_{4}$ of the PQG state, only terms with $\mathcal{D}$ acting outermost survive, which yields
@@
For simplicity, we only consider the nearest-neighbor terms $V$ that dominate in $V_{ij}\propto|i-j|^{-6}$. In the small-$\Delta_{l}$ limit, the first-order $\mathcal{O}(\Delta_{l})$ contribution in Eq. (S16b) yields density-dressed single-bubble and pair-exchange blocks,
@@
When evaluating the expectation value of $\hat{C}_{4}$ of the Néel state, the lowest term in the $\Delta_{l}$-polynomial is the zeroth-order, $\mathcal{O}\left((\Delta_{l})^{0}\right)$. Using the definition of the Néel state, Eq. (S16a) yields:
@@ IV.D Higher orders and practical truncation
At higher orders, in the small-$\Delta_{l}$ limit one still finds $\langle\hat{C}_{2m}\rangle_{\mathrm{PQG}}=\mathcal{O}(\Delta_{l})$, indicating the slow decay of PQG state. At order $\hat{C}_{6}$ we encounter length-3 (tri-bubble) NN strings, e.g.